What would be the change in volume if 60 ml of a gas is cooled from 33.0 0C to 5.00 0C?


1
Expert's answer
2022-03-23T23:13:02-0400
  • We will use the formula V1/T1=V2/T2
  • Note that temperature should be in kelvin so we add 273.15
  • 60/306.15=V2/278.15
  • This gives 54.51ml